OpenStudy (anonymous):

solve this exponential equation 8^x-3 = 16

OpenStudy (anonymous):

is ur question 8^(x-3) = 16

OpenStudy (anonymous):

yes solve for x

OpenStudy (anonymous):

(2^3)^(x-3) = 2^4 or 2^(3*(x-3) =2^4 or 3(x-3) = 4 ...and so on

OpenStudy (anonymous):

thank you =)